What is the equation of a line graphed through (-2,0)(0,-2)
Answer by checkley71(8403)

(-2-0)/(0-2)
-2/2=-1
Y=-X+b SUBSTITUTING ONE SET OFPOINTS WE FIND THE Y INTERCEPT
0=-1*-2+b
0=2+b
b=-2THUS THE EQUATION IS
Y=-X-2